OpenStudy (anonymous):

1.) [4^3 sqrt. a] + [^3 sqrt. 64a] 2.) [sqrt. 5] - [7 sqrt. 80] + [4 sqrt. 45]

OpenStudy (anonymous):

u guyz ought to help everyone.......plz help me also.......

OpenStudy (anonymous):

If you can convert the bit under the radical into a product of some squares you can pull out the square and you might end up with something you can combine with addition/subtraction.

OpenStudy (anonymous):

Sidd, please stop stalking me.

OpenStudy (anonymous):

2) -15 Sqrt[5] not exactly what the first is asking though

OpenStudy (anonymous):

siddharth tiwari : ....... well i need help with these so ? ......

OpenStudy (anonymous):

a933793: please don't just give people answers. If they want the solution there's wolfram alpha. If they want to learn it'll take more explanation than what you've provided.

OpenStudy (anonymous):

k..........just got anxious...u kno....when u dont have any way out to do the quesn ,,,,how does it feels.......well dont mind sir....was just a request not an order.....sry if it mean here

OpenStudy (anonymous):

Kathy, can you factor out some of the product under the sqrt and make it into a product (multiplication) of some squared factors?

OpenStudy (anonymous):

i have no idea polopak , i get so confused with these sqrts :/

OpenStudy (anonymous):

Don't worry about the square roots for now. Can you factor 80?

OpenStudy (anonymous):

well the common gcf of all 5, 80, 45 is 5

OpenStudy (anonymous):

for problem 2

OpenStudy (anonymous):

Not greatest common factor, Just rewrite them as a product of their factors. \(80 = 8 *10 = (4*2) * (5*2) \) \(= (2*2*2)*(5*2) = (2*2)*(2*2) * 5\) \(= 2^2 * 2^2 * 5\)

OpenStudy (anonymous):

Now try doing the same thing for 45.

OpenStudy (anonymous):

And be sure to group up any squares you find

OpenStudy (anonymous):

1,5,9,45 = factors for 45

OpenStudy (anonymous):

Ignore 1 and itself.

OpenStudy (anonymous):

and for 5 , it would be just 5

OpenStudy (anonymous):

45 = 9*5, and 9 is what?

OpenStudy (anonymous):

3*3

OpenStudy (anonymous):

\(3^2\)

OpenStudy (anonymous):

Ok, so lets rewrite the problem now with the squares.. \(\sqrt{5} - 7\sqrt{2^2 * 2^2 * 5} + 4\sqrt{3^2 * 5}\)

OpenStudy (anonymous):

Now any square factor we can pull out of the square root by removing the square.

OpenStudy (anonymous):

\(7*\sqrt{2^2 * 2^2 * 5} = 7 * \sqrt{2^2} * \sqrt{2^2} * \sqrt{5} = ?\)

OpenStudy (anonymous):

i multiply the sqrts ?

OpenStudy (anonymous):

No, take the square root of the squared term. What is \(\sqrt{2^2}\)

OpenStudy (anonymous):

If a >= 0, then \(\sqrt{a^2} = a\)

OpenStudy (anonymous):

Are you still with me?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, great. So what's \(\sqrt{2^2}\)

OpenStudy (anonymous):

sqrt. 4

OpenStudy (anonymous):

Yes, but what's sqrt of 4? It's 2 right? You don't need to evaluate \(2^2\) because when you take the square root you just take off the square and the square root. They undo eachother. \(\sqrt{2^2} = 2. \) \(\sqrt{3^3} = 3.\) etc.

OpenStudy (anonymous):

Err \(\sqrt{3^2} = 3\)

OpenStudy (anonymous):

so it would just be 2 and 3 ?

OpenStudy (anonymous):

Not quite. So we have \(7*\sqrt{80} = 7*\sqrt{2^2 *2^2 * 5} = 7*\sqrt{2^2}*\sqrt{2^2} * \sqrt{5} =?\)

OpenStudy (anonymous):

7 * 2 sqrt. 5 ?

OpenStudy (anonymous):

correction : 7 * 2 * 2 * sqrt. 5 ?

OpenStudy (anonymous):

Right. So now simplify that

OpenStudy (anonymous):

7*2*2 = ?

OpenStudy (anonymous):

becuase their are 2 [sqrt.2^2]

OpenStudy (anonymous):

No, we got rid of the square root already on those. \((7*2*2)\sqrt{5} \) \(= (14*2 ) \sqrt{5}\) \(= 28 \sqrt{5}\)

OpenStudy (anonymous):

can we still simplfy that ?

OpenStudy (anonymous):

No, but we have \(\sqrt{5} - 28\sqrt{5} + 4\sqrt{45}\) So if we can get the \(4\sqrt{45}\) to be a multiple of \(\sqrt{5}\) we can combine them all.

OpenStudy (anonymous):

So lets look at \(\sqrt{45}\). We said that that was 45 = 9*5 = \(3^2*5\), so what would \(4*\sqrt{45}\) simplify to?

OpenStudy (anonymous):

2 sqrt. 5 ?

OpenStudy (anonymous):

4 sqrt . 5 *3^2

OpenStudy (anonymous):

So what is the next step now that you have that?

OpenStudy (anonymous):

Just like the last one, we pull out each squared factor.

OpenStudy (anonymous):

4*3 sqrt.5 ?

OpenStudy (anonymous):

Yes. Which is?

OpenStudy (anonymous):

What's 4*3?

OpenStudy (anonymous):

12 sqrt. 5

OpenStudy (anonymous):

Right. So now we have \(1\sqrt{5} - 28\sqrt{5} + 12\sqrt{5}\)

OpenStudy (anonymous):

-15 sqrt. 5 ?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

and for the 1st problem , i factor out which is a ?

OpenStudy (anonymous):

Do the same thing. Factor everything under the square root. Pull out any square factors by removing the square, leave the rest under the square root. Then combine.

OpenStudy (anonymous):

Go ahead and do it yourself, step by step and I'll point out when/if you get off track.

OpenStudy (anonymous):

Just post each step as you go.

OpenStudy (anonymous):

What's the first step?

OpenStudy (anonymous):

[4^3 sqrt. a] + [^3 sqrt. 64a] : what do i do with the powers before the sqrt ?

OpenStudy (anonymous):

are you still here ?

OpenStudy (anonymous):

I don't understand.. Do you have cube roots? \(4\sqrt[3]{a} + \sqrt[3]{64a}\) If so then instead of squares, you're looking for cubes. \(\sqrt[3]{2^3} = 2 \) is an example you will probably find useful. If you are still working on this, give it a try and write out your steps and where you're stuck and I'll check back again later.

OpenStudy (anonymous):

yes i still need help please

OpenStudy (anonymous):

polpak: i think it'll result to 16\[^{3} \sqrt{a}\]

OpenStudy (anonymous):

What was the first step you did?